show me how to solve 6/2(2+1)=



Answer by Shana-D77(132)

PEMDAS
(), exponent, mult/divide, add/subtract.
6/2(2+1)=
6/2(3)=
Remember, 3 is the same as 3/1, so:
(6/2)(3/1)=
always with fractions, multiply numerator*numerator and denominator*denominator:
18/2=
9